Find mean, median of the following data
CI f
20-24 8
25-29 16
30-34 20
35-39 18
40-44 10
45-49 7
50-54 3
55-59 2

Answer by Boreal(15235)

the mean is the midpoint of each range multiplied by the number and all added
22.5*8=176, because it is 20-24.99
27.5*16=440
32.5*20=650
37.5*18=675
42.5*10=425
47.5*7=332.5
52.5*3=156.75
57.5*2=115
sum is 2970.25
divide by n=84
mean=35.36
median is midway between 42nd and 43rd value and is 32.5, since it is all within the 30-34.9 range.